For science fair I need to calculate the area of mold on bread. I was told to use a grid. So I made a 1/2 cm by 1/2 cm grid. Now I don't know how to use the grid to find the area.could someone tell me how to use a 1/2 cm by 1/2 cm grid to find the area of mold?
Answer by Alan3354(69443)

Area = 1/2 * 1/2 = 1/4 sq cm for each square of the grid